OpenStudy (anonymous):

Number of years 1 2 3 Option 1 (amount in dollars) 1300 1690 2197 Option 2 (amount in dollars) 1300 1600 1900 Belinda wants to invest in an option that would help to increase her investment value by the greatest amount in 20 years. Will there be any significant difference in the value of Belinda's investment after 20 years if she uses option 2 over option 1? Explain your answer, and show the investment value after 20 years for each option.

OpenStudy (danjs):

you cn see from the first 3 values for years 1,2,and 3, Option 1 increases in value faster than Option 2

OpenStudy (danjs):

Option 2 is linear , with a slope of 300 dollars per year Option 1 is exponential

OpenStudy (danjs):

For Option 1, use the general form y = a*b^x

OpenStudy (danjs):

use a point (x,y) = (1,1300) 1300 = a*b^1 Another point (x1,y1) = (2,1690) 1690 = a*b^2 2 equations, solvable for constants a and b

OpenStudy (danjs):

does that form y = a*b^x look familiar from class ?

OpenStudy (anonymous):

no not really

OpenStudy (danjs):

Ok, lets start with the easier one, option 2.....do you see that each year the value increases by the same amount, 300 dollars?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

Since the value increases at the same amount each year, the graph will be a line, we can use 2 of the points from the table to find the slope of the line. In general, the line will be in the form: y = m*x where m is the slope.

OpenStudy (danjs):

using points (1,1300) and (2,1600) the slope m, is the difference in y over differene in x m = (1600-1300)/(2-1) = 300/1 = 300 The line has a slope of 300 dollars per year.

OpenStudy (anonymous):

oh ok

OpenStudy (danjs):

and starts out at year 1 with 1300 dollars, so the equation of the line is , using the point- slope formula for a line, point (1,1300) and slope m=300 y - 1300 = 300(x-1) y - 1300 = 300x - 300

OpenStudy (danjs):

The equation for OPTION 2 is: \[\large\color{blue}{ y = 300x + 1000}\]

OpenStudy (danjs):

where x is the year, and y is the money value

OpenStudy (anonymous):

ohh okayy thanks so much

OpenStudy (danjs):

After 20 years, OPTION 2 will be worth y = 300(20) + 1000 \[\large\color{blue}{OPTION~2 ~at ~20~years = $7000 }\]

OpenStudy (danjs):

That was for the linear option 2 For the OPTION 1, we have to find the Exponential equation

OpenStudy (anonymous):

okay

OpenStudy (danjs):

The general form for an exponential equation is \[\large\color{blue}{ y = C*a^x}\] we need to use points from the table to figure out what C and a will be..you follow?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

Here are 2 points from OPTION 1: (x,y) (1,1300) and (2,1690)

OpenStudy (danjs):

Lets use the first point in the general form above (x,y) = (1,1300) \[1300 = C*a^1\] Now use the second point in the same general equation: (x,y) = (2,1690) \[1690 = C*a^2\] see that ?

OpenStudy (danjs):

We have 2 equations, with 2 unknown variables 'C' and 'a' \[\large\color{blue}{1300 = C*a^1~~~and~~~1690 = C*a^2 }\]

OpenStudy (danjs):

Solving the first equation for C gives, \[C = \frac{ 1300 }{ a }\] Putting that C value into the second equation gives, \[1690 = C*a^2 = \frac{ 1300 }{ a }a^2\]

OpenStudy (danjs):

Now you can solve for what 'a' is: 1690 = (1300/a)*a^2 1690 = 1300a a = 13/10

OpenStudy (danjs):

Now that 'a' is known , we can find C... 1300 = C*(13/10)^1 1300*10 = 13C C= 1000

OpenStudy (danjs):

Overall the exponential equation for the OPTION 1 is given by: \[y = 1000*(\frac{ 13 }{ 10 })^x\] where y is the value in dollars, and x is the time in years

OpenStudy (danjs):

As a test, let x=3 for the third year: y = 1000*(13/10)^3 = 2197 That is true and the equation works, at year 3, the value is 2197 dollars

OpenStudy (danjs):

So at year 20, (x=20), the value of OPTION 1 is: y = 1000*(13/10)^20 = 190050 dollars

OpenStudy (danjs):

at the 20 year Mark: OPTION 1 = $190,050 OPTION 2 = $7000 Big difference.